An unopened can of soda has an aqueous CO2 concentration of 0.21 M at 25.0
Answer #1

K = C/P , 3.4 x 10^-2 = 0.21 /P

P = 6.17 atm = 6.2 atm = 6.2 x 760 torr = 4.7 x 10^3 torr

hence A is answer